OpenStudy (anonymous):

ln(2n+3) = ln 3 Solve the Equation?

zepdrix (zepdrix):

\[\large \ln(2n+3)=\ln3\] We'll start by exponentiating both sides (rewrite both sides as exponents with a base of e). \[\huge e^{\ln(2n+3)}=e^{\ln3}\] The exponential function and the logarithmic function are inverses. So they essentially undo one another. Giving us,\[\large 2n+3=3\]

zepdrix (zepdrix):

If the exponent thing was confusing.. maybe ignore that. Just realize that since we're taking the log of both sides, we can set the contents of the logs equal to one another.

OpenStudy (anonymous):

In other words, e to the natural logarithm of 2n+3 is just 2n+3, and e to the natural logarithm of 3 is 3.

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

youre welcome. Like 10 to the power of log n is n. :)